# LLMs as Bridges: Reformulating Grounded Multimodal Named Entity Recognition

### v3

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) is a nascent multimodal task that aims to identify named entities, entity types and their corresponding visual regions. GMNER task exhibits two challenging properties: 1) The weak correlation between image-text pairs in social media results in a significant portion of named entities being ungroundable. 2) There exists a distinction between coarse-grained referring expressions commonly used in similar tasks (e.g., phrase localization, referring expression comprehension) and fine-grained named entities. In this paper, we propose RiVEG, a unified framework that reformulates GMNER into a joint MNER-VE-VG task by leveraging large language models (LLMs) as a connecting bridge. This reformulation brings two benefits: 1) It maintains the optimal MNER performance and eliminates the need for employing object detection methods to pre-extract regional features, thereby naturally addressing two major limitations of existing GMNER methods. 2) The introduction of entity expansion expression and Visual Entailment (VE) Module unifies Visual Grounding (VG) and Entity Grounding (EG). It enables RiVEG to effortlessly inherit the Visual Entailment and Visual Grounding capabilities of any current or prospective multimodal pretraining models. Extensive experiments demonstrate that RiVEG outperforms state-of-the-art methods on the existing GMNER dataset and achieves absolute leads of 10.65%, 6.21%, and 8.83% in all three subtasks.
@@ 1 Introduction
Multimodal Named Entity Recognition (MNER) (Lu et al., 2018a) aims to extract named entities and corresponding categories from image-text pairs sourced from social media. As one of the most important subtask in multimodal information processing, MNER has received extensive attention in recent years (Wang et al., 2022c, b; Li et al., 2023). Due to the pressing demand for constructing structured knowledge graphs, human-machine interaction and robotics, recent research endeavors to extend the scope of MNER (Wang et al., 2023a). As a further extension, Grounded Multimodal Named Entity Recognition (GMNER) (Yu et al., 2023) aims to extract text named entities, entity types and their bounding box groundings from image-text pairs.
@@
Analyzing GMNER from the two perspectives of MNER and EG, existing GMNER methods (Yu et al., 2023) have several significant limitations. Limitation 1. Entity Grounding (EG) is an important part in GMNER, which aims to determine the groundability and visual grounding results of text named entities predicted by the MNER model. Obviously, effective EG depends on accurate MNER. But the MNER performance of the existing multimodal fusion GMNER methods (Yu et al., 2023) is suboptimal. This may be due to the introduction of visual object features destroying the original text feature representation in MNER (Wang et al., 2022c). A series of MNER studies in the Text-Text paradigm (i.e., represent images using text) (Wang et al., 2022b; Li et al., 2023) show that text representation plays the most critical role in MNER. Models with the best MNER performance typically do not involve any cross-modal interactions. However, existing end-to-end GMNER methods (Yu et al., 2023) violate this in order to obtain EG capabilities with the help of cross-modal interactions. This results in the feature representation learned by the model being neither the best for MNER nor the best for EG. Therefore, splitting the prediction and grounding of named entities into separate stages is necessary. Limitation 2. Existing GMNER methods (Yu et al., 2023; Wang et al., 2023a) typically employ object detection (OD) techniques to extract features from candidate regions, which are then aligned with text entities. But as shown in Table 1, our statistical results111Here, Acc@0.5 refers to the accuracy metric in object detection across all images. It measures the proportion of images where at least one of the Top-N predicted bounding boxes based on detection probability has an IoU of 0.5 or greater with the ground-truth bounding box. This proportion is then expressed as the accuracy rate. indicate that the widely-adopted OD methods exhibit unsatisfactory performance in recognizing candidate visual objects in GMNER dataset. This implies that there exists a natural upper limit to the effectiveness of existing GMNER methods. In instances where the OD methods are unable to detect the visual ground-truth bounding boxes for textual named entities, even if the MNER part predicts the correct textual named entity, the EG result will definitely be wrong. Therefore, it is necessary to avoid extracting visual features from candidate image regions.
@@
Visual Grounding (VG) (Rohrbach et al., 2016; Yu et al., 2018) aims to correlate natural language descriptions with specific regions in images. Although VG has a similar paradigm to the Entity Grounding (EG) in GMNER, existing VG methods (Zhu et al., 2022; Deng et al., 2021; Wang et al., 2022a) exhibit significant deficiencies when applied to the GMNER dataset. This is attributed to two fundamental distinctions between these two tasks: Differentiation 1. Referring to the Pipeline-1 and Pipeline-2 in Figure 1, the text input for the EG task consists of real-world named entities, which is notably different from the referring expressions or noun phrases typically used in the VG method. There is a substantial difference that is almost non-generalizable between them. The VG method can comprehend the description "A female superhero in the MCU" but struggles with understanding named entities like "Black Widow". Therefore, a bridge between noun phrases and named entities needs to be built.
@@
Differentiation 2. Due to the social media usage patterns of users, image-text pairs from social media exhibit apparent weak correlation. It implies that the textual named entity may not necessarily be linked to a specific region in the image. Yu et al. (2023) shows that approximately 60% of the named entities in the GMNER dataset are ungroundable. But the classic VG task stipulates that the input referring expression must match an object in the image, without considering ungroundable expressions that do not match any object. It means that the behavior of the existing VG methods is undefined if the named entity does not correspond to any object in the image. As shown in Pipeline-1 and Pipeline-2 of Figure 1, existing VG methods inevitably make errors when they confront ungroundable text input like "Marvel Studios". Our experiments demonstrate that the Acc@0.5 of the state-of-the-art VG methods applied to the GMNER dataset is 21.87% (Wang et al., 2022a) and 9.23% (Zhu et al., 2022), respectively.222We fine-tune VG models using the GMNER dataset. Text inputs are named entities and entity types. For ungroundable entities, the coordinates of gold labels are defined as all zeros. Therefore, a separate module is needed to determine the groundability of text inputs.
@@
RiVEG demonstrates a new paradigm for GMNER, increases the application scope of the VG model to the cases of no right objects in image. This paradigm supports data augmentation and facilitates subsequent upgrades.
@@ 2.1 Multimodal Named Entity Recognition
MNER is a relatively independent multimodal subtask. Early MNER methods attempt to simply interact text and images (Moon et al., 2018; Lu et al., 2018b; Yu et al., 2020). To deal with the limitations of this shallow interaction, subsequent methods explore various cross-modal fusion approaches (Zhang et al., 2021a; Wang et al., 2022e; Jia et al., 2023). There is a trend to solve MNER through the Text-Text paradigm (Wang et al., 2022c, b; Li et al., 2023), and this paradigm generally yields better performance. Despite obtaining promising results, these methods lack any visual understanding capabilities. The GMNER task poses significant challenges to the above research.
@@ 2.2 Grounded Multimodal Named Entity Recognition and Visual Grounding
GMNER is a further exploration of MNER. Existing GMNER works (Yu et al., 2023; Wang et al., 2023b) aim to utilize OD methods for obtaining region proposals and subsequently sorting them based on their relevance to named entities. Such methods are inevitably limited by OD methods. VG is a representative multimodal subtask. The progress in multimodal pretraining research (Zhu et al., 2022; Wang et al., 2022a; Yan et al., 2023) continues to contribute to the improvement of VG methods. However, the VG task has some strong predefined constraints. It presupposes that there are objects in the image that match the input text query. This renders VG methods inapplicable to the GMNER task that deviates from such predefined constraints. In this paper, we introduce the VE module to reconcile this conflict. Our method uses the VG method as the main body of visual understanding and achieves better performance than the existing GMNER methods.
@@ 3 Methodology
RiVEG is mainly divided into two stages. In the stage of Named Entity Recognition and Expansion, we employ auxiliary refined knowledge obtained from LLMs to ensure optimal MNER performance. Furthermore, in order to adapt existing VG methods to EG task, LLMs is considered as a bridge between named entities and referring expressions. RiVEG instructs LLMs to convert named entities into named entity referring expressions (detailed in §3.2). In the stage of Named Entity Grounding, RiVEG reformulates the entire EG task as a union of VE and VG. To reconcile the challenge between the weak correlation of image-text pairs in GMNER and the strong predefined constraints of existing VG methods, we design the VE module for buffering. Images and named entity referring expressions serve as the input to VE module for determining the groundability of the named entity. Finally, the named entities deemed groundable alongside images are employed as inputs to VG module to obtain the final grounding results (detailed in §3.3). We also design a data augmentation method suitable for RiVEG. An overview of the RiVEG is depicted in Figure 2.
@@ 3.1 Task Formulation
Given a sentence with tokens and its corresponding image , the objective of GMNER is to identify and categorize named entities (i.e., , , and ) within sentences while concurrently determining the visually grounded region corresponding to each entity. Formulate that the output of GMNER comprises a set of multimodal entity triples:
@@
Incorporating document-level extended knowledge into original multimodal samples has been proven to be one of the most effective methods for MNER (Wang et al., 2022c, b; Li et al., 2023). PGIM (Li et al., 2023) points out that despite the inherent limitations of generative models in sequence labeling tasks, LLMs serve as an implicit knowledge base that is highly suitable for MNER task. We use the auxiliary refined knowledge333Original text and image captions are used as input to LLMs, and the LLMs are guided to generate extended explanations of the original samples through in-context learning. provided by PGIM to enhance the predictive performance of named entities. And we additionally use the same algorithm to generate different auxiliary refined knowledge for the same sample from more versions of LLMs to perform data augmentation. Specifically, definite auxiliary refined knowledge is retrieved by LLMs as . We take the concatenation of the original sentence together with the auxiliary refined knowledge as the input of the transformer-based encoder:
@@
where is the potential function and are predefined semantic categories with the BIO2 tagging schema (Sang and Veenstra, 1999). Finally, given the input sequence with gold labels , we employ the negative log-likelihood (NLL) loss function for training model:
@@
where is reserved for LLMs to generate. A limited number of predefined in-context examples are employed to guide LLMs in response format (detailed in Appendix Figure 5). Finally, we concatenate named entities, entity categories, and entity expansion expressions into the final named entity referring expressions:
@@
For existing VG methods, fine-tuning with limited data hardly equips them with the ability to effectively handle irrelevant text and images. Hence, addressing the weak correlation between image and text in GMNER poses the primary challenge in integrating VG methods into GMNER. To achieve this objective, RiVEG introduces the VE module before the VG module.
@@
where , , represent a named entity corresponding image, a named entity referring expression and a class label. Note that since the groundability of a named entity in GMNER dataset is unambiguous, only two labels and are retained. represents is groundable and represents is ungroundable. Finally, RiVEG replaces the original VE dataset with to fine-tune the existing vanilla VE model. In the inference phase, only the samples deemed groundable by the fine-tuned vanilla VE model will be fed into the final VG module. For the remaining samples that are judged to be ungroundable, their entity grounding results are directly defined as .
Visual Grounding Module
Expansion expressions bring named entities closer to the referring expressions required by existing VG methods. And VE module effectively filters out ungroundable samples. Once there is no requirement for VG methods to assess the groundability of text input, existing vanilla VG methods can be seamlessly applied to determine the visually grounded region of the named entity. Denote the subset of all groundable named entity samples in GMNER dataset as:
where , , represent a named entity corresponding image, a named entity referring expression and a visually grounded region of , respectively. Note that if a named entity in corresponds to multiple different bounding boxes, we specify that the bounding box with the largest area is chosen as the unique gold label. Subsequently, RiVEG employs to substitute the original VG dataset and fine-tune the existing vanilla VG model.
4 Experiments
@@ 4.1 Settings
Based on two benchmark MNER datasets, i.e., Twitter-2015 (Zhang et al., 2018) and Twitter-2017 (Yu et al., 2020), Yu et al. (2023) builds the Twitter-GMNER dataset by filtering samples with missing images or more than 3 entities belonging to the same type. More details about the training data for different modules and the Twitter-GMNER dataset are provided in the Appendix A.2.
@@ 4.2 Main Results
Note that this result does not represent the performance upper limit of RiVEG. The modular design of RiVEG facilitates straightforward upgrades, and selecting more advanced methods for different modules in the future is expected to yield improved performance.444Furthermore, we test 10 other textual and visual variations of RiVEG in detail in Appendix 10 and A.6. Experimental results show that the performance of all 14 RiVEG variants is superior to the previous state-of-the-art method.
@@ 4.3 Detailed Analysis
| LLMs in GMNER(Acc@0.5) | SeqTR | OFA |
@@
Table 3 shows the test set results after using the same MNER model trained on different LLM versions of the same dataset. The method of obtaining external knowledge is exactly the same as PGIM (Li et al., 2023). Compared with ITA (Wang et al., 2022c) (Using optical character recognition, object detection and image caption) and MoRe (Wang et al., 2022b) (Using Wikipedia), the auxiliary knowledge provided by LLMs is more helpful to MNER. The quality of knowledge provided by LLMs with a small number of parameters is usually inferior to ChatGPT used by PGIM. RiVEG performs better than past state-of-the-art methods.555This is average result from three different random seeds. The training set of RiVEG includes responses from all five LLMs. And the validation set and test set are the same as PGIM. This indicates that utilizing diverse styles of external knowledge based on the same original sample enhances the generalization ability of model.
The named entities and their visual grounding results in the Twitter-GMNER dataset are deterministic, so using different LLMs to generate different named entity expansion expressions can build different LLM versions of the dataset. Table 4 shows the test set results of the same VE and VG models after training on different versions of the datasets. The training set of the Mix version is a direct merger of all five training sets, and the validation set is identical to the ChatGPT version. Using all versions of the Named Entity Referring Expression as text input performs significantly better than using raw named entities. LLMs with a large number of parameters generally generate better quality results than LLMs with a small number of parameters. The results of the Mix version show that more different styles of generated data can increase the generalization of the model. This is reflected more intuitively in Appendix A.6.
@@
| VG Module Text input (Acc@0.5) | SeqTR | OFA |
@@
We further investigate the performance of various methods in VE and VG modules when confronted with diverse text inputs. The named entity referring expressions in are segmented into the named entity along with entity category and the entity expansion expression. Table 5 shows the test set performance of different methods after fine-tuning on . In order to control variables, this experiment is conducted on the expansion expression generated by ChatGPT.
@@
As the datasets we create differs only in text type from the SNLI-VE (Xie et al., 2019) and RefCOCO series datasets (Mao et al., 2016; Yu et al., 2016; Nagaraja et al., 2016), we investigate whether pretraining with more similar task data could yield additional performance improvements in Table 6.
For various VE methods, incorporating additional fine-tuning on the SNLI-VE dataset (Xie et al., 2019) prior to the final fine-tuning on the does not result in improved performance. This observation suggests a significant difference between the and the SNLI-VE. For various VG methods, conducting additional fine-tuning on the RefCOCO series of datasets (Mao et al., 2016; Yu et al., 2016; Nagaraja et al., 2016) prior to the final fine-tuning on the may result in enhanced performance. This may be attributed to a certain degree of similarity between the two. Because we convert named entities into named entity referring expressions through entity expansion expressions.666ALBEF (Li et al., 2021) authors do not provide pre-trained weights based on the SNLI-VE dataset. And some of the pre-trained weights of SeqTR (Zhu et al., 2022) based on the RefCOCO series dataset cannot be obtained. Therefore, we follow the wishes of author and only test OFA in this experiment.

### v4

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) is a nascent multimodal task that aims to identify named entities, entity types and their corresponding visual regions. GMNER task exhibits two challenging properties: 1) The weak correlation between image-text pairs in social media results in a significant portion of named entities being ungroundable. 2) There exists a distinction between coarse-grained referring expressions commonly used in similar tasks (e.g., phrase localization, referring expression comprehension) and fine-grained named entities. In this paper, we propose RiVEG, a unified framework that reformulates GMNER into a joint MNER-VE-VG task by leveraging large language models (LLMs) as a connecting bridge. This reformulation brings two benefits: 1) It maintains the optimal MNER performance and eliminates the need for employing object detection methods to pre-extract regional features, thereby naturally addressing two major limitations of existing GMNER methods. 2) The introduction of entity expansion expression and Visual Entailment (VE) module unifies Visual Grounding (VG) and Entity Grounding (EG). It enables RiVEG to effortlessly inherit the Visual Entailment and Visual Grounding capabilities of any current or prospective multimodal pretraining models. Extensive experiments demonstrate that RiVEG outperforms state-of-the-art methods on the existing GMNER dataset and achieves absolute leads of 10.65%, 6.21%, and 8.83% in all three subtasks. Code and datasets publicly available at https://github.com/JinYuanLi0012/RiVEG.
@@ 1 Introduction
Multimodal Named Entity Recognition (MNER) (Lu et al., 2018a) aims to extract named entities and corresponding categories from image-text pairs sourced from social media. As one of the most important subtask in multimodal information processing, MNER has received extensive attention in recent years (Wang et al., 2022c, b; Li et al., 2023; Cui et al., 2024). Due to the pressing demand for constructing structured knowledge graphs, human-machine interaction and robotics, recent research endeavors to extend the scope of MNER (Wang et al., 2023). As a further extension, Grounded Multimodal Named Entity Recognition (GMNER) (Yu et al., 2023) aims to extract text named entities, entity types and their bounding box groundings from image-text pairs.
@@
Analyzing GMNER from the two perspectives of MNER and EG, existing GMNER methods (Yu et al., 2023) have several significant limitations. Limitation 1. Entity Grounding (EG) is an important part in GMNER, which aims to determine the groundability and visual grounding results of text named entities predicted by the MNER model. Obviously, effective EG depends on accurate MNER. But the MNER performance of the existing multimodal fusion GMNER methods (Yu et al., 2023) is suboptimal. This may be due to the introduction of visual object features destroying the original text feature representation in MNER (Wang et al., 2022c). A series of MNER studies in the Text-Text paradigm (i.e., represent images using text) (Wang et al., 2022b; Li et al., 2023) show that text representation plays the most critical role in MNER. Models with the best MNER performance typically do not involve any cross-modal interactions. However, existing end-to-end GMNER methods (Yu et al., 2023) violate this in order to obtain EG capabilities with the help of cross-modal interactions. This results in the feature representation learned by the model being neither the best for MNER nor the best for EG. Therefore, splitting the prediction and grounding of named entities into separate stages is necessary. Limitation 2. Existing GMNER methods (Yu et al., 2023; Wang et al., 2023) typically employ object detection (OD) techniques to extract features from candidate regions, which are then aligned with text entities. But as shown in Table 1, our statistical results111Here, TopN-Prec@0.5 refers to the accuracy metric in object detection methods across all samples. It measures the proportion of named entities where at least one of the Top-N predicted bounding boxes based on detection probability has an IoU of 0.5 or greater with the ground-truth bounding box. indicate that the widely-adopted OD methods exhibit unsatisfactory performance in recognizing candidate visual objects in GMNER dataset. This implies that there exists a natural upper limit to the effectiveness of existing GMNER methods. In instances where the OD methods are unable to detect the visual ground-truth bounding boxes for textual named entities, even if the MNER part predicts the correct textual named entity, the EG result will definitely be wrong. Therefore, it is necessary to avoid extracting visual features from candidate image regions.
@@
Visual Grounding (VG) (Rohrbach et al., 2016; Yu et al., 2018) aims to correlate natural language descriptions with specific regions in images. Although VG has a similar paradigm to the Entity Grounding (EG) in GMNER, existing VG methods (Zhu et al., 2022; Deng et al., 2021; Wang et al., 2022a) exhibit significant deficiencies when applied to the GMNER dataset. This is attributed to two fundamental distinctions between these two tasks: Differentiation 1. Referring to the Pipeline-1 and Pipeline-2 in Figure 1, the text input for the EG task consists of real-world named entities, which is notably different from the referring expressions or noun phrases typically used in the VG method. There is a substantial difference that is almost non-generalizable between them. The VG method can comprehend the description “A female superhero in the MCU” but struggles with understanding named entities like “Black Widow”. Therefore, a bridge between noun phrases and named entities needs to be built.
@@
Differentiation 2. Due to the social media usage patterns of users, image-text pairs from social media exhibit apparent weak correlation. It implies that the textual named entity may not necessarily be linked to a specific region in the image. Yu et al. (2023) shows that approximately 60% of the named entities in the GMNER dataset are ungroundable. But the classic VG task stipulates that the input referring expression must match an object in the image, without considering ungroundable expressions that do not match any object. It means that the behavior of the existing VG methods is undefined if the named entity does not correspond to any object in the image. As shown in Pipeline-1 and Pipeline-2 of Figure 1, existing VG methods inevitably make errors when they confront the ungroundable text input like “Marvel Studios”. Our experiments demonstrate that the Prec@0.5 of the state-of-the-art VG methods applied to the GMNER dataset is 21.87% (Wang et al., 2022a) and 9.23% (Zhu et al., 2022), respectively.222We fine-tune VG models using the GMNER dataset. Text inputs are named entities and entity types. For ungroundable entities, the coordinates of gold labels are defined as all zeros. Therefore, a separate module is needed to determine the groundability of text inputs.
@@
RiVEG increases the application scope of VG models to the cases of no right objects in image and demonstrates a new paradigm for GMNER. This paradigm supports data augmentation and facilitates subsequent upgrades.
@@ 2.1 Multimodal Named Entity Recognition
MNER is a relatively independent multimodal subtask. Early MNER methods attempt to simply interact text and images (Moon et al., 2018; Lu et al., 2018b; Yu et al., 2020). To deal with the limitations of this shallow interaction, subsequent methods explore various cross-modal fusion approaches (Zhang et al., 2021a; Wang et al., 2022e; Jia et al., 2023; Chen et al., 2022a, b). There is a trend to solve MNER through the Text-Text paradigm (Wang et al., 2022c, b; Li et al., 2023), and this paradigm generally yields better performance. Despite obtaining promising results, these methods lack any visual understanding capabilities. The GMNER task poses significant challenges to the above research.
@@ 2.2 Grounded Multimodal Named Entity Recognition and Visual Grounding
GMNER is a further exploration of MNER. Existing GMNER works (Yu et al., 2023; Wang et al., 2023) aim to utilize OD methods for obtaining region proposals and subsequently sorting them based on their relevance to named entities. Such methods are inevitably limited by OD methods. Visual Grounding (VG) is a representative multimodal subtask. The progress in multimodal pretraining research (Zhu et al., 2022; Wang et al., 2022a; Yan et al., 2023) continues to contribute to the improvement of VG methods. However, the VG task has some strong predefined constraints. It presupposes that there are objects in the image that match the input text query. This renders VG methods inapplicable to the GMNER task that deviates from such predefined constraints. In this paper, we introduce the VE module to reconcile this conflict. Our method uses the VG method as the main body of visual understanding and achieves better performance than the existing GMNER methods.
@@ 3 Methodology
RiVEG is mainly divided into two stages. In the stage of Named Entity Recognition and Expansion, we employ auxiliary refined knowledge obtained from LLMs to ensure optimal MNER performance. Furthermore, in order to adapt existing VG methods to EG task, LLMs is considered as a bridge between named entities and referring expressions. RiVEG instructs LLMs to convert named entities into named entity referring expressions (detailed in §3.2). In the stage of Named Entity Grounding, RiVEG reformulates the entire EG task as a union of VE and VG. To reconcile the challenge between the weak correlation of image-text pairs in GMNER and the strong predefined constraints of existing VG methods, we design the VE module for buffering. Images and named entity referring expressions serve as the input to VE module for determining the groundability of the named entities. Finally, the named entities deemed groundable alongside images are employed as inputs to VG module to obtain the final grounding results (detailed in §3.3). We also design a data augmentation method suitable for RiVEG. An overview of the RiVEG is depicted in Figure 2.
@@ 3.1 Task Formulation
Given a sentence with tokens and its corresponding image , the objective of GMNER is to identify and categorize named entities (i.e., , , and ) within the sentence while concurrently determining the visually grounded region corresponding to each entity. Formulate that the output of GMNER comprises a set of multimodal entity triples:
@@
Incorporating document-level extended knowledge into original multimodal samples has been proven to be one of the most effective methods for MNER (Wang et al., 2022c, b; Li et al., 2023). PGIM (Li et al., 2023) points out that despite the inherent limitations of generative models in sequence labeling tasks, LLM serve as an implicit knowledge base that is highly suitable for MNER task. We use the auxiliary refined knowledge333Original text and image captions are used as inputs to LLMs, and LLMs are guided to generate extended explanations of the original samples. provided by PGIM to enhance the predictive performance of named entities. And we additionally use the same algorithm to generate different auxiliary refined knowledge for the same sample from more versions of LLMs to perform data augmentation. Specifically, definite auxiliary refined knowledge retrieved by LLMs as . We take the concatenation of the original sentence together with the auxiliary refined knowledge as the input of the transformer-based encoder:
@@
where is the potential function and is the set of all possible label sequences given the input and . Finally, given the input sequence with gold labels , we employ the negative log-likelihood (NLL) loss function for training model:
@@
where is reserved for LLMs to generate. A limited number of predefined in-context examples are employed to guide the format of responses (detailed in Appendix Figure 5). Finally, we concatenate named entities, entity categories, and entity expansion expressions into the final named entity referring expressions:
@@
For existing VG methods, fine-tuning with limited data hardly equips them with the ability to effectively handle irrelevant text and images. Hence, addressing the weak correlation between images and text in GMNER poses the primary challenge in integrating VG methods into GMNER. To achieve this objective, RiVEG introduces the VE module before the VG module.
@@
where , , represent an image that corresponds to a named entity, a named entity referring expression and a class label, respectively. Note that since the groundability of a named entity in GMNER dataset is unambiguous, only two labels and are retained. represents is groundable and represents is ungroundable. Finally, RiVEG replaces the original VE dataset with to fine-tune the existing vanilla VE model. In the inference phase, only the samples deemed groundable by the fine-tuned vanilla VE model will be fed into the final VG module. For the remaining samples that are judged to be ungroundable, their entity grounding results are directly defined as .
@@
Visual Grounding Module
Expansion expressions bring named entities closer to the referring expressions required by existing VG methods. And VE module effectively filters out ungroundable samples. Once there is no requirement for VG methods to assess the groundability of the text input, existing vanilla VG methods can be seamlessly applied to determine the visually grounded region of the named entity. Denote the subset of all groundable named entity samples in GMNER dataset as:
where , , represent an image that corresponds to a named entity, a named entity referring expression and a visually grounded region of , respectively. Note that if a named entity in corresponds to multiple different bounding boxes, we specify that the bounding box with the largest area is chosen as the unique gold label. Subsequently, RiVEG employs to substitute the original VG dataset and fine-tune the existing vanilla VG model.
4 Experiments
@@ 4.1 Settings
Based on two benchmark MNER datasets, i.e., Twitter-2015 (Zhang et al., 2018) and Twitter-2017 (Yu et al., 2020), Yu et al. (2023) builds the Twitter-GMNER dataset by filtering samples with missing images or more than 3 entities belonging to the same type. More details about the training data for different modules and the Twitter-GMNER dataset are provided in Appendix A.2.
@@ 4.2 Main Results
Note that this result does not represent the performance upper limit of RiVEG. The modular design facilitates straightforward upgrades, selecting more advanced methods for different modules in the future is expected to yield improved performance.444Furthermore, we test 10 other text and visual variations of RiVEG in detail in Appendix 10 and A.6. Experimental results show that the performance of all 14 RiVEG variants is superior to the previous state-of-the-art method.
@@ 4.3 Detailed Analysis
| LLMs in GMNER(Prec@0.5) | SeqTR | OFA |
@@
Table 3 shows the test set results after using the same MNER model trained on different LLM versions of the same dataset. The method of obtaining external knowledge is exactly the same as PGIM (Li et al., 2023). Compared with ITA (Wang et al., 2022c) (Using optical character recognition, object detection and image caption) and MoRe (Wang et al., 2022b) (Using Wikipedia), the auxiliary knowledge provided by LLMs is more helpful to MNER. The quality of knowledge provided by LLMs with a small number of parameters is usually inferior to ChatGPT used by PGIM. RiVEG performs better than past state-of-the-art methods.555This is average result from three different random seeds. The training set of RiVEG includes responses from all five LLMs. And the dev set and test set are the same as PGIM. This indicates that utilizing diverse styles of external knowledge based on the same original sample enhances the generalization ability of model.
The named entities and their visual grounding results in the Twitter-GMNER dataset are deterministic, so using different LLMs to generate different named entity expansion expressions can build different LLM versions of the dataset. Table 4 shows the test set results of the same VE and VG models after training on different versions of the datasets. The training set of the Mix version is a direct merger of all five training sets, and the dev set is identical to the ChatGPT version. Using all versions of the named entity referring expression as text input performs significantly better than using raw named entities. LLMs with a large number of parameters generally generate better quality results than LLMs with a small number of parameters. The results of the Mix version show that more different styles of generated data can improve model performance. This is reflected more intuitively in Appendix A.6.
@@
| VG Module Text Input (Prec@0.5) | SeqTR | OFA |
@@
We further investigate the performance of various methods in VE and VG modules when confronted with diverse text inputs. The named entity referring expression in is segmented into the named entity along with entity category and the entity expansion expression. Table 5 shows the test set performance of different methods after fine-tuning on . In order to control variables, this experiment is conducted on the expansion expression generated by ChatGPT.
@@
As the datasets we create differ only in text type from the SNLI-VE (Xie et al., 2019) and RefCOCO series datasets (Mao et al., 2016; Yu et al., 2016; Nagaraja et al., 2016), we investigate whether pretraining with more similar task data could yield additional performance improvements in Table 6.
For various VE methods, incorporating additional fine-tuning on the SNLI-VE dataset (Xie et al., 2019) prior to the final fine-tuning on the does not result in improved performance. This observation suggests a significant difference between the and the SNLI-VE. For various VG methods, conducting additional fine-tuning on the RefCOCO series of datasets (Mao et al., 2016; Yu et al., 2016; Nagaraja et al., 2016) prior to the final fine-tuning on the may result in enhanced performance. This may be attributed to a certain degree of similarity between the two. Because we convert named entities into named entity referring expressions through entity expansion expressions.666ALBEF (Li et al., 2021) authors do not provide pretrained weights based on the SNLI-VE dataset. And some of the pretrained weights of SeqTR (Zhu et al., 2022) based on the RefCOCO series dataset cannot be obtained. Therefore, we follow the wishes of author and only test OFA(Wang et al., 2022a) in this experiment.